A flat circle of radius 18 cm is placed in a uniform electric field of magnitude 5.8×102 N/C. What is the electric flux through the circle when its face is: 


1. perpendicular to the field lines

2. 45° to the field lines and

3. parallel to the field lines?



1
Expert's answer
2021-02-24T16:58:41-0500
"\Phi=EA\cos{\alpha}"

1)


"\Phi=\pi(0.18)^2(580)\cos{0}=59\ Vm"

2)


"\Phi=\pi(0.18)^2(580)\cos{45}=42\ Vm"

3)


"\Phi=\pi(0.18)^2(580)\cos{90}=0\ Vm"